# ToMCAT: Theory-of-Mind for Cooperative Agents in Teams via Multiagent Diffusion Policies

### v1

@@ 3.2 Theory-of-Mind Reasoning
As mentioned earlier, our first goal is to create a mechanism capable of inferring the motivations and intent of teammates from their observed behavior. To achieve that, we follow the approach in (Rabinowitz et al., 2018) which aims to develop a machine ToM by training a neural network called a ToMnet to model the mental states and behaviors of other agents through meta-learning. The ToMnet learns to predict agents’ future actions and infer their goals and beliefs, including false beliefs, by observing their behavior in an environment of interest.
@@
The ToMnet is parametrized by $\theta$, $\phi$ and $\gamma$ and is trained end-to-end. It leverages meta-learning to construct a general ToM (the model’s parameters encode a prior for modeling diverse agent types) while dynamically updating an agent-specific ToM (the model’s predictions update a posterior over the characteristics of a particular teammate from recently observed behavior (Rabinowitz et al., 2018).Our extension to multiagent settings results in a single model capable of forming predictions about multiple agent types from the perspective of different types of observer agents.
@@ 4       Agent $i$ observes $o_{i}$; $\texttt{enqueue}(h,o_{i})$;
6      if $\textnormal{empty}(\tau_{i})\ ||\ \lVert\eta(o_{i})-\eta(\hat{o_{i}})\rVert^{2}>\lambda$ then
@@ 20       Execute $a_{i}$, let teammates $j\in N^{-{i}}$ act;
As such, we propose a dynamic replanning approach where an agent follows the actions in a generated plan while its predictions, including about its teammates, remain consistent, and replans whenever they are inconsistent or the plan is depleted. Moreover, to reduce the number of denoising steps, we use DDIM sampling (Song et al., 2021). This results in the online planning (sampling) approach delineated in Algorithm 1, corresponding to a modified version of the algorithm in (Ajay et al., 2023). In addition to computing and using $e_{\text{char}}$, $e_{\text{ment}}$ and $\psi_{i}$ as conditioning variables, the other novelty resides in line 4, where we compute the MSE between the current observation and the observation predicted in the plan, where $\eta:\Omega\to[0,1]$ is an observation normalizer. If the observation difference is higher than a predefined threshold, $\lambda$, we generate a new plan, otherwise continue using the actions from the previously computed one. See Sec. A.1 of the Appendix for a detailed description of the algorithm.
@@ 4.4.2 Known Teammates
In this experiment we want to address the question: “Can ToM reasoning improve team performance in the presence of strong teammate priors?” Here we want to understand whether ToM is important for ToMCAT agents when paired with teammates for whom prior behavior demonstrations are provided, corresponding to supplying $N_{\text{past}}=4$ past trajectories as detailed in the previous experiment. To address that question, we test ToMCAT agents equipped with MADiff models under different conditioning variables, $y({\tau_{i}})$, namely: Unconditioned, Returns ($[R_{\text{task}}]$), Profile ($[\psi_{i}]$), and ToM ($[\psi_{i},e_{\text{char}},e_{\text{ment}}]$). As an upper bound on performance, we also test against a team of RL agents, where at each episode we sample a pair of policies that were concurrently trained via MARL, corresponding to the Know RL condition.
@@ 4.4.3 Unknown Teammates
Table 3 shows the results for the different conditions. Unsurprisingly, the pure RL team attained the best mean individual performance since their policies were trained to maximize the reward function associated with their profiles; however, they failed to coordinate with their teammate due to the random policy selection. We also see that despite not being provided prior information about a teammate, the ToM agent is still able to perform better than the unconditioned agent and the RL team, although the difference was not statistically significant ($p\gg 0.05$). Moreover, the Profile condition attained the highest mean task performance.
@@ A.1 Dynamic Replanning Algorithm
Our dynamic replanning algorithm (Alg. 1) unfolds as follows: for a given agent $i$, we first initialize the character embeddings for teammates from past trajectories, if available. We also initialize empty queues for the observation history, $h$, and the planned sequence of observations, $\tau_{\text{plan}}$. At each step, after the agent makes a new observation in the environment, we check if replanning is needed (line 4). If a plan has been previously generated and is not depleted, we compute the mean squared error between the current observation and the observation predicted in the plan, where $\eta:\Omega\to[0,1]$ is an observation normalizer, which is detailed below in Sec. A.5.2. If the observation difference is higher than a predefined threshold, $\lambda$, we generate a new plan (lines 5–11). First, the mental embedding is computed from the current trajectory and all conditioning variables—any combination of returns ($R_{i}(\tau_{i})$), observer profile ($\psi_{i}$), character embedding ($e_{\text{char}}$), mental embedding ($e_{\text{ment}}$)—are concatenated into $y({\tau_{i}})$ (line 5). We then generate a new plan through conditional sampling, starting with $\hat{\boldsymbol{\tau}}^{K}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. At each diffusion step $k$, the portion of the observer agent in $\hat{\boldsymbol{\tau}}^{k}$ is “in-painted” with the corresponding current observation and $C$ steps of observation history (line 8). We iteratively refine (denoise) trajectories $\hat{\boldsymbol{\tau}}^{k}$ using the perturbed noise $\hat{\epsilon}$ following (Zhu et al., 2024; Ajay et al., 2023), where $\epsilon_{\delta}(\hat{\boldsymbol{\tau}}^{k},k)$ is the unconditional noise, $\omega$ is a scalar for extracting the distinct portions of the trajectory that exhibit $y({\tau_{i}})$, and denoise is the noise scheduler predicting the previous sample, $\hat{\boldsymbol{\tau}}^{k-1}$, from $\hat{\boldsymbol{\tau}}^{k}$ by propagating the diffusion process from $\hat{\epsilon}$. Once the final, denoised trajectory, $\hat{\boldsymbol{\tau}}^{0}$, is generated, we discard the first $C+1$ steps—since they were only used to condition the generation process—and extract agent $i$’s portion, $\hat{\tau}_{i}$, re-initializing the plan queue, $\tau_{\text{plan}}$ (line 11). The agent then retrieves an action by using the inverse dynamics model $I_{\xi}$ on consecutive observations extracted from the plan (line 12). [^7] Here the dequeue function extracts observations from the queue in a FIFO manner.
@@ A.5.2 Training
For each sampled trajectory (length $H+C$) we augmented the observer agent’s observations by including the one-hot encoding of the teammate’s actions and then performed Cumulative Density Function (CDF) based normalization, corresponding to $\eta$ in Alg. 1. For returns conditioning, we computed the discounted cumulative task reward in the original trajectory starting from the timestep in which the MADiff trajectory was sampled and then divided it by the maximum task reward that agents can receive, i.e.20202020 in our case corresponding to a soup delivery. The reason for considering the whole episode instead of just the forward horizon for computing returns is that this way we denote the future potential of the sampled trajectory in achieving the task’s goal (delivering soups), and not just what happens within the sample window itself. For the other conditioning variables, i.e., observer profile, and character and mental embeddings, we normalized them using CDF similarly to observations. In summary, a single datapoint used for training the diffusion model consisted of a window of augmented observations, a return, an observer profile vector, and a character and a mental embedding.

### v2

@@ 3.2 Theory-of-Mind Reasoning
As mentioned earlier, our first goal is to create a mechanism capable of inferring the motivations and intent of teammates from their observed behavior. To achieve that, we follow the approach in (Rabinowitz et al., 2018), which aims to develop a machine ToM by training a neural network called a ToMnet to model the mental states and behaviors of other agents through meta-learning. The ToMnet learns to predict agents’ future actions and infer their goals and beliefs, including false beliefs, by observing their behavior in an environment of interest.
@@
The ToMnet is parametrized by $\theta$, $\phi$ and $\gamma$ and is trained end-to-end. It leverages meta-learning to construct a general ToM—the model’s parameters encode a prior for modeling diverse agent types—while dynamically updating an agent-specific ToM — the model’s predictions update a posterior over the characteristics of a particular teammate from recently observed behavior Rabinowitz et al. (2018). Our extension to multiagent settings results in a single model capable of forming predictions about multiple agent types from the perspective of different types of observer agents. Since others’ behavior is contingent on an observer’s own behavior, the observer’s profile is used as input to condition the predictions of the different modules.
@@ 4       Agent $i$ observes $o_{i}$; $\texttt{enqueue}(h,o_{i})$;
6      if $\textnormal{empty}(\tau_{i})\ ||\ \delta_{\text{obs}}(o_{i},\hat{o_{i}})>\lambda$ then
@@ 20       Execute $a_{i}$, let teammates $j\in N^{-{i}}$ act;
As such, we propose a dynamic replanning approach where an agent follows the actions in a generated plan while its predictions, including about its teammates, remain consistent, and replans whenever they are inconsistent or the plan is depleted. Moreover, to reduce the number of denoising steps, we use DDIM sampling (Song et al., 2021). This results in the online planning (sampling) approach delineated in Algorithm 1, corresponding to a modified version of the algorithm in (Ajay et al., 2023). In addition to computing and using $e_{\text{char}}$, $e_{\text{ment}}$ and $\psi_{i}$ as conditioning variables, the other novelty resides in line 4, where we compute the difference between the current observation and the observation predicted in the plan, using a suitable metric, $\delta_{\text{obs}}$. If the observation difference is higher than a predefined threshold, $\lambda$, we generate a new plan, otherwise we continue using the actions from the previously computed one. See Sec. A.1 of the Appendix for a detailed description of the algorithm.
@@ 4.4.2 Known Teammates
In this experiment, we want to address the question: “Can ToM reasoning improve team performance in the presence of strong teammate priors?” Here we want to understand whether ToM is important for ToMCAT agents when paired with teammates for whom prior behavior demonstrations are provided, corresponding to supplying $N_{\text{past}}=4$ past trajectories as detailed in the previous experiment. To address that question, we test ToMCAT agents equipped with MADiff models under different conditioning variables, $y({\tau_{i}})$, namely: Unconditioned, Returns ($[R_{\text{task}}]$), Profile ($[\psi_{i}]$), and ToM ($[\psi_{i},e_{\text{char}},e_{\text{ment}}]$). As an upper bound on performance, we also test against a team of RL agents, where at each episode we sample a pair of policies that were concurrently trained via MARL, corresponding to the Know RL condition.
@@ 4.4.3 Unknown Teammates
Table 3 shows the results for the different conditions. The pure RL agents failed to coordinate with their teammates due to the random policy selection. Notwithstanding, they attained the best mean individual performance since their policies were trained to maximize the reward function associated with their profiles, irrespectively of the teammate. We also see that despite not being provided prior information about a teammate, the ToM agent is still able to perform better than the unconditioned agent and the RL team (difference not statistically significant, $p\gg 0.05$). Moreover, the Profile condition attained the highest mean task performance.
@@ A.1 Dynamic Replanning Algorithm
Our dynamic replanning algorithm (Alg. 1) unfolds as follows: for a given agent $i$, we first initialize the character embeddings for teammates from past trajectories, if available. We also initialize empty queues for the observation history, $h$, and the planned sequence of observations, $\tau_{\text{plan}}$.
At each step, after the agent makes a new observation in the environment, we check if replanning is needed (line 4). If a plan has been previously generated and is not depleted, we compute the mean squared error between the current observation and the observation predicted in the plan, i.e., we set $\delta_{\text{obs}}\coloneq\lVert\eta(o_{i})-\eta(\hat{o_{i}})\rVert^{2}$, where $\eta:\Omega\to[0,1]$ is an observation normalizer detailed in Sec. A.5.2.
If the observation difference is greater than $\lambda=0.2$, we generate a new plan (lines 5–11). First, the mental embedding is computed from the current trajectory, and all conditioning variables—any combination of returns ($R_{i}(\tau_{i})$), observer profile ($\psi_{i}$), character embedding ($e_{\text{char}}$), mental embedding ($e_{\text{ment}}$)—are concatenated into $y({\tau_{i}})$ (line 5). We then generate a new plan through conditional sampling, starting with $\hat{\boldsymbol{\tau}}^{K}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. At each diffusion step $k$, the portion of the observer agent in $\hat{\boldsymbol{\tau}}^{k}$ is “in-painted” with the corresponding current observation and $C$ steps of observation history (line 8). We iteratively refine (denoise) trajectories $\hat{\boldsymbol{\tau}}^{k}$ using the perturbed noise $\hat{\epsilon}$ following Zhu et al. (2024); Ajay et al. (2023), where $\epsilon_{\delta}(\hat{\boldsymbol{\tau}}^{k},k)$ is the unconditional noise, $\omega$ is a scalar for extracting the distinct portions of the trajectory that exhibit $y({\tau_{i}})$, and denoise is the noise scheduler predicting the previous sample, $\hat{\boldsymbol{\tau}}^{k-1}$, from $\hat{\boldsymbol{\tau}}^{k}$ by propagating the diffusion process from $\hat{\epsilon}$.
Once the final, denoised trajectory, $\hat{\boldsymbol{\tau}}^{0}$, is generated, we discard the first $C+1$ steps—since they were only used to condition the generation process—and extract agent $i$’s portion, $\hat{\tau}_{i}$, re-initializing the plan queue, $\tau_{\text{plan}}$ (line 11). The agent then retrieves an action by using the inverse dynamics model $I_{\xi}$ on consecutive observations extracted from the plan (line 12). [^7] Here the dequeue function extracts observations from the queue in a FIFO manner.
@@ A.5.2 Training
For each sampled trajectory (length $H+C$), we augmented the observer agent’s observations by including the one-hot encoding of the teammate’s actions and then performed Cumulative Density Function (CDF) based normalization, corresponding to function $\eta$ in Alg. 1. For returns conditioning, we computed the discounted cumulative task reward in the original trajectory starting from the timestep in which the MADiff trajectory was sampled and then divided it by the maximum task reward that agents can receive, i.e.20202020 in our case corresponding to a soup delivery. The reason for considering the whole episode instead of just the forward horizon for computing returns is that this way we denote the future potential of the sampled trajectory in achieving the task’s goal (delivering soups), and not just what happens within the sample window itself. For the other conditioning variables, i.e., observer profile, and character and mental embeddings, we normalized them using CDF similarly to observations. In summary, a single datapoint used for training the diffusion model consisted of a window of augmented observations, a return, an observer profile vector, and a character and a mental embedding.